in an isosceles right triangle the hypotenuse is 12 inches how long are the other legs?
Answer by unlockmath(1688)

Hello,
Since we know the other sides are the same we can set up the following equation:
x^2+x^2=12^2
Rewritten as:
2x^2=144
Divide by 2 and square root both sides to get:
x=sq rt 72
There we go. So the lengths of both sides are sq rt 72.
Make sense?
RJ
www.math-unlock.com